OpenStudy (anonymous):

Ann is driving down a street at 59 km/h. Suddenly a child runs into the street. If it takes Ann 0.765 s to react and apply the brakes, how far will she have moved before she begins to slow down? Answer in units of m

OpenStudy (anonymous):

Start by converting the 59 km/h to km/s.

OpenStudy (anonymous):

How do you do that? @gypsy1274

OpenStudy (anonymous):

How many minutes in an hour? How many seconds in a minute?

OpenStudy (anonymous):

60 minutes in an hour and 60 seconds in a minute @gypsy1274

OpenStudy (anonymous):

1 hour \(\times\) 60 minutes \(\times\) 60 seconds

OpenStudy (anonymous):

FYI: you don't need to tag me, since I have already posted a reply, I will get a notice when anyone posts another reply.

OpenStudy (anonymous):

oh sorry and its 3,600

OpenStudy (anonymous):

Now set up a proportion. The speed per second on one side.

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

\(\dfrac{km}{seconds} = \dfrac{km}{seconds}\) The first fraction could be the driving speed. What you don't know is how far she will travel - that can be represented by x.

OpenStudy (anonymous):

would it be \[\frac{ 59 }{ 0.765 }=\frac{ x }{ 0.765 }\]

OpenStudy (anonymous):

Almost. Remember that 59km is per hour or 3600 seconds. Make that correction, then cross multiply.

OpenStudy (anonymous):

\[\frac{ 3600 }{ 0.765 }=\frac{ x }{ 0.765 } answer is 36000\]

OpenStudy (anonymous):

@gypsy1274

OpenStudy (anonymous):

@chmvijay can you please help me?

OpenStudy (chmvijay):

57 km/h/3600 s/h=.0158333 km/s=15.83 m/s 15.83 m/s x .765 s= 12.1125 m..

OpenStudy (anonymous):

the answer is 12.1125m/s? @CHM

OpenStudy (anonymous):

@chmvijay

OpenStudy (anonymous):

@Hero is the answer 12.1125m/s

OpenStudy (anonymous):

that was the wrong answer @chmvijay

OpenStudy (anonymous):

Online question for hw

OpenStudy (chmvijay):

how do u assure that i solved one is wrong ?

OpenStudy (anonymous):

because it said incorrect @ch

OpenStudy (chmvijay):

ok ok fine

OpenStudy (anonymous):

would the answer be 3600m

OpenStudy (chmvijay):

14000

OpenStudy (anonymous):

i just crossed multiplied\[\frac{ 3600 }{ 0.765 }\times \frac{ x }{ 0.765 }\]

OpenStudy (anonymous):

answer 3600 is that right?

OpenStudy (anonymous):

@chmvijay

OpenStudy (chmvijay):

IDK

OpenStudy (chmvijay):

@satellite73 @dumbcow @kropot72 @robtobey @radar @timo86m @dan815 @Luigi0210 cn u answer this question :)

OpenStudy (dumbcow):

@chmvijay , you did it correctly but you misread the speed ....its 59 km/hr not 57 :)

OpenStudy (chmvijay):

but though it wont make so much different in answer right right and she said "my answer is showing wrong in online test"

OpenStudy (abb0t):

0.8 seconds? That baby dead...

OpenStudy (dumbcow):

haha someone was texting and driving

OpenStudy (dumbcow):

@JazminZamarripa \[\frac{59 km}{hr}*\frac{1 hr}{3600 \sec}*\frac{1000 m}{1 km} = \frac{16.388 m}{\sec}\] multiply by .765 sec to gave distance in meters car travels

OpenStudy (anonymous):

Thank you so much:) @dumbcow

OpenStudy (chmvijay):

no @abb0t that baby is still alive :P

OpenStudy (dumbcow):

yw

OpenStudy (chmvijay):

hey thank you @dumbcow

OpenStudy (anonymous):

Thank you too @chmvijay

OpenStudy (chmvijay):

yw :)